# arXiv:2602.14013v2[math.CV] 01 Mar 2026

### v2

@@ 2 Preliminaries
Due to [MRSY, Section 3] we say that a boundary $D$ is called strongly accessible with respect to $p$-modulus at $x_{0}\in\partial D,$ if for any neighborhood $U$ of the point $x_{0}\in\partial D$ there is a neighborhood $V\subset U$ of this point, a compactum $F\subset D$ and a number $\delta>0$ such that $M_{p}(\Gamma(E,F,D))\geqslant\delta$ for any continua $E\subset D$ such that $E\cap\partial U\neq\varnothing\neq E\cap\partial V.$ The boundary of a domain $D$ is called strongly accessible with respect to $p$-modulus, if this is true for any $x_{0}\in\partial D.$ When $p=n,$ prefix ’’relative to $p$-modulus’’ is omitted.
Remark 2.1. Uniform domains have strongly accessible boundaries, see e.g. [SevSkv, Remark 1]. The inverse statement is also true at least for $p=n,$ see [Sev3, Lemma 4.2].
The following statement is true, see [Sev2, Lemma 2.1].
Lemma 2.1.  Let $p\geqslant 1,$ $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ be an open discrete ring $Q$-mapping at the point $b\in\partial D$ with respect to the $p$-modulus, $b\neq\infty,$ $f(D)=D^{\,\prime}$, a domain $D$ is locally connected at the point $b,$ $C(f,\partial D)\subset\partial D^{\,\prime}$ and the domain $D^{\,\prime}$ is strongly accessible with respect to the $p$-modulus at least at one point $y\in C(f,b),$ where $C(f,b)$ is a cluster set of $f$ at $b.$ Suppose that there is $\varepsilon_{0}>0$ and some positive measurable function $\psi:(0,\varepsilon_{0})\rightarrow(0,\infty)$ such that
$0<I(\varepsilon,\varepsilon_{0})=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt<\infty$ (2.1)
@@
$\int\limits_{A(\varepsilon,\varepsilon_{0},b)}Q(x)\cdot\psi^{\,p}(|x-b|)\ dm(x)=o(I^{p}(\varepsilon,\varepsilon_{0}))\,$, (2.2)
@@
Given $p\geqslant 1,$ a number $\delta>0,$ a family $\mathfrak{D}$ of domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D})$ a family of all homeomorphisms $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ defined in some $D\in\mathfrak{D}$ and satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{\mathfrak{D}}$ for which there exists a continuum $A=A_{f}\subset D$ with $h(f(A))\geqslant\delta.$
@@
as $\varepsilon\rightarrow 0,$ where $A(x_{0},\varepsilon,\varepsilon_{0})$ is defined in (1.2).
@@
Proof.   We will mainly use the methodology that has been used for mappings of a fixed domain, see, e.g., [SevSkv]. Assume the contrary. Then there exists $a>0$ such that for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in B(x_{0},1/k)\cap D_{m_{k}}$ such that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ We may consider that the sequence $m_{k}$ is increasing by $k.$ Otherwise, $m_{k}=k_{0}$ for sufficiently large $k\in{\mathbb{N}}$ and some $k_{0}\in{\mathbb{N}},$ besides that $x_{0}\in\overline{D_{k_{0}}}$ because $f_{k_{0}}$ is defined at $x_{k}\cap D_{k_{0}}$ and $x_{k}\rightarrow x_{0}$ as $k\rightarrow\infty.$ Now, from the assumption above we have that $h(f_{k_{0}}(x_{k}),f_{k_{0}}(x^{\,\prime}_{k}))\geqslant a.$ The latter contradicts with Lemma 2. So, $m_{k}$ may be chosen as an increasing sequence. Without loss of generality, going to renumbering, if required, we may consider that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a$ holds for $m$ instead of $m_{k},$ i.e.,
@@
where $h$ denotes the chordal diameter of the set.
Let $\Gamma_{m}$ be a family of paths joining $\gamma_{m}$ and $A_{m}$ in $D_{m},$ where $A_{m}$ is a continuum from the condition 3). By the definition of the class $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D}),$ the relations (1.3)–(1.4) hold at $x_{0}.$ Therefore,
$M_{p}(f_{m}(\Gamma_{m}))\leqslant\int\limits_{A(x_{0},\frac{1}{2^{m}},\varepsilon_{0})}Q(x)\cdot\eta^{p}(|x-x_{0}|)\ dm(x)$ (2.7)
for every Lebesgue measurable function $\eta:(\frac{1}{2^{m}},\varepsilon_{0})\rightarrow[0,\infty]$ such that $\int\limits_{\frac{1}{2^{m}}}^{\varepsilon_{0}}\eta(r)\,dr\geqslant 1.$ Observe that, the function
@@
The following statement may be found in [Sev2, Lemmas 1.3, 1.4].
@@
$\int\limits_{\delta}^{\delta_{0}}\frac{dt}{t^{\frac{n-1}{p-1}}q_{x_{0}}^{\frac{1}{p-1}}(t)}<\infty,\qquad 0<\delta<\delta_{0},$ (2.9)
and
$\int\limits_{0}^{\delta_{0}}\frac{dt}{t^{\frac{n-1}{p-1}}q_{x_{0}}^{\frac{1}{p-1}}(t)}=\infty\,.$ (2.10)
Then there exist a number $\varepsilon_{0}\in(0,1)$ and a function $\psi(t)\geqslant 0$ satisfying (2.3) such that the relation (2.4) holds.
@@ 3 On image of the continuum under open discrete and closed mappings
Proposition 3.2.   Let $f:D\rightarrow{\mathbb{R}}^{n}$ be a discrete open and closed (boundary preserving) mapping, $\beta:[a,\,b)\rightarrow f(D)$ be a path, and $x\in\,f^{\,-1}\left(\beta(a)\right).$ Then $\beta$ has a total $f$-lifting starting at $x$ (see, e.g., [Vu, Lemma 3.7]).
@@
Lemma 3.2.  Let $D_{0}\subset{\mathbb{R}}^{n},$ be a domain in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and let $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ be a sequence of domains the kernel of which is $D_{0}.$ Let $A$ be a non-degenerate continuum $A\subset D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ and let $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$ be a Lebesgue measurable function. In addition, let $f_{m}:D_{m}\rightarrow{\mathbb{R}}$ be a sequence of open discrete and closed mappings satisfying relations (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ for all $0<r_{1}<r_{2}<d(x_{0},\partial D_{m})$ such that $h(f_{m}(A))\geqslant\delta>0$ for any $m=1,2,\ldots$ and $p=n$ for some $\delta>0.$ Assume that, $Q$ has a finite mean oscillation in $A,$ or the relation (1.6) holds for every $x_{0}\in A$ and some $\beta(x_{0})>0.$ If there is $r>0$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m}),$ then there exists $\delta_{1}>0$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$
Proof.   Observe that, $h(A,\partial D_{m})\geqslant r_{*}>0$ for some $r_{*}>0,$ and all $m\in{\mathbb{N}}$ except finite numbers. Otherwise, there are sequences $x_{m}\in A$ and $y_{m}\in\partial D_{m}$ with $h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty.$ We may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ Obviously, $x_{0}\in A$ by the closeness of $A.$ Now, there is $\varepsilon_{0}>0$ such that $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}$ and $\overline{B(x_{0},\varepsilon_{0})}$ belongs to all $D_{m}$ except of finite numbers. However, the latter contradicts the condition $y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty$ while $y_{m}\in\partial D_{m}.$ Indeed, in this case $\overline{B(x_{0},\varepsilon_{0})}\cap\partial D_{m}\neq\varnothing$ for infinitely many $m\in{\mathbb{N}}.$
@@
the relation (1.6) implies that, given $0<\widetilde{\varepsilon_{1}}<r_{*}\leqslant h(A,\partial D_{m})$ we may find $\varepsilon_{1}=\varepsilon_{1}(x_{0})$ such that
$\int\limits_{\varepsilon_{1}}^{\widetilde{\varepsilon_{1}}}\frac{dt}{tq_{x_{0}}^{\,\prime\,\frac{1}{n-1}}(t)}>1\,.$ (3.1)
We cover $A$ with the balls $B(x_{0},\varepsilon_{1}),$ $x_{0}\in A,$ where $\varepsilon_{1}$ as above. By Heine-Borel-Lebesgue lemma we may find a finite subcover $B(x_{1},\varepsilon_{1}),$ $B(x_{2},\varepsilon_{2}),\ldots,B(x_{N},\varepsilon_{N})$ such that
$A\subset\bigcup\limits_{i=1}^{n}B(x_{i},\varepsilon_{i})\,,\qquad A\subset\bigcup\limits_{i=1}^{n}B(x_{i},2\varepsilon_{i})\,.$ (3.2)
By the choice of $\varepsilon_{i}>0,$ $\varepsilon_{i}<\min\{\beta(x_{0}),\frac{1}{2}r_{*}\},$ we have that $B(x_{i},2\varepsilon_{i})\subset D_{m}.$ Since $D_{m}$ is a domain in ${\mathbb{R}}^{n},$ $m=1,2,\ldots,$ $\partial D_{m}\neq\varnothing.$ In addition, since $C(\partial D_{m},f_{m})\subset\partial f_{m}(D_{m})$ whenever $f_{m}$ is open, discrete and closed (see [Vu, Theorem 3.3]), $\partial f_{m}(D_{m})\neq\varnothing,$ as well. Thus, the quantity $h(f_{m}(A),\partial f_{m}(D_{m}))$ is well-defined, so the formulation of the lemma is correct.
@@
Put $P>0$ and $U=B_{h}(y_{0},r_{0})=\{y\in\overline{{\mathbb{R}}^{n}}:h(y,y_{0})<r_{0}\},$ where $2r_{0}:=\min\{r/2,\delta/2\},$ where $r$ and $\delta$ are numbers from the condition of the lemma. Observe that $E_{m}\cap U\neq\varnothing\neq E_{m}\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $y_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $y_{m}\in E_{m};$ besides that $h(E_{m})\geqslant r>r/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Observe that, $E_{m}$ is closed (see [Ku, Theorem 1.III.5.46]). Now, $E_{m}$ is a continuum, so that $E_{m}\cap\partial U\neq\varnothing$ by Proposition 3. Similarly, $f_{m}(A)\cap U\neq\varnothing\neq f_{m}(A)\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $x_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $x_{m}\in f_{m}(A);$ besides that $h(f_{m}(A))\geqslant\delta>\delta/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Since $f_{m}(A)$ is a continuum, $f_{m}(A)\cap\partial U\neq\varnothing$ by Proposition 3. By the proving above,
$E_{m}\cap\partial U\neq\varnothing\neq f_{m}(A)\cap\partial U$ (3.3)
for sufficiently large $m\in{\mathbb{N}}.$ By Lemma 3 there is $\subset U, VV$ is a neighborhood of $y_{0},$ such that
@@
for any continua $E,F\subset\overline{{\mathbb{R}}^{n}}$ with $E\cap\partial U\neq\varnothing\neq E\cap\partial V$ and $F\cap\partial U\neq\varnothing\neq F\cap\partial V.$ Arguing similarly to above, we may prove that
$E_{m}\cap\partial V\neq\varnothing\neq f_{m}(A)\cap\partial V$
for sufficiently large $m\in{\mathbb{N}}.$ Thus, by (3.4)
@@
Let $\Delta_{m}$ be a family of all total $f_{m}$ liftings of $\Gamma_{m}$ starting at $A$ (it exists by Proposition 3). Now
@@
Let $\beta_{m}:[0,t_{m})\rightarrow D_{m},$ $\beta_{m}\in\Delta_{m},$ $f_{m}\circ\beta_{m}=\alpha_{m}.$ Since open discrete and closed mappings preserve the boundary ([Vu, Theorem 3.3]), $\beta_{m}(t)\rightarrow\partial D_{m}$ as $t\rightarrow t_{m}.$
@@
Indeed, let $\gamma\in\Delta_{m},$ in other words, $\gamma:[0,t_{m})\rightarrow D_{m},$ $\gamma(0)\in A$ and $\gamma(t)\rightarrow\partial D_{m}$ as $t\rightarrow t_{m}-0.$ Now, there is $1\leqslant i\leqslant N$ such that $|\gamma|\cap B(x_{i},\varepsilon_{i})\neq\varnothing\neq|\gamma|\cap(D_{m}\setminus B(x_{i},\varepsilon_{i})).$ Therefore, by Proposition 3 there is $0<t^{\,*}_{1}<t_{m}$ such that $\gamma(t^{\,*}_{1})\in S(x_{i},\varepsilon_{i}).$ We may assume that $\gamma(t)\not\in B(x_{i},\varepsilon_{i})$ for $t>t^{\,*}_{1}.$ Put $\gamma_{1}:=\gamma|_{[t^{\,*}_{1},t_{m}]}.$ Similarly, $|\gamma_{1}|\cap B(x_{i},2\varepsilon_{i})\neq\varnothing\neq|\gamma_{1}|\cap(D_{m}\setminus B(x_{i},2\varepsilon_{i})).$ By Proposition 3 there is $t^{\,*}_{1}<t^{\,*}_{2}<t_{m}$ with $\gamma(t^{\,*}_{2})\in S(x_{i},2\varepsilon_{i}).$ We may assume that $\gamma(t)\in B(x_{i},\varepsilon_{i})$ for $t<t^{\,*}_{2}.$ Put $\gamma_{2}:=\gamma|_{[t^{\,*}_{1},t^{\,*}_{2}]}.$ Then, the path $\gamma_{2}$ is a subpath of $\gamma,$ which belongs to the family $\Gamma(S(x_{i},\varepsilon_{i}),S(x_{i},2\varepsilon_{i}),A(x_{i},\varepsilon_{i},2\varepsilon_{i})).$ Thus, the relation (3.8) is established.

### v3

@@ 2 Preliminaries
The following statement is true, see [Sev1, Lemma 2.1].
Lemma 2.1.  Let $p\geqslant 1,$ and let $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ be an open discrete ring $Q$-mapping at the point $b\in\partial D$ with respect to $p$-modulus, $b\neq\infty,$ $f(D)=D^{\,\prime}$, a domain $D$ is locally connected at the point $b,$ $C(f,\partial D)\subset\partial D^{\,\prime}$ and $D^{\,\prime}$ is strongly accessible with respect to the $p$-modulus at least at one point $y\in C(f,b),$ where $C(f,b)$ is a cluster set of $f$ at $b.$ Suppose that there is $\varepsilon_{0}>0$ and some positive measurable function $\psi:(0,\varepsilon_{0})\rightarrow(0,\infty)$ such that
$0<I(\varepsilon,\varepsilon_{0})=\int\limits_{\varepsilon}^{\varepsilon_{0}}\psi(t)\,dt<\infty$
@@
$\int\limits_{A(b,\varepsilon,\varepsilon_{0})}Q(x)\cdot\psi^{\,p}(|x-b|)\ dm(x)=o(I^{p}(\varepsilon,\varepsilon_{0}))\,$,
@@
Given $p\geqslant 1,$ a number $\delta>0,$ a family $\mathfrak{D}$ of domains in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and a Lebesgue measurable function $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$, we denote by $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D})$ a family of all homeomorphisms $f:D\rightarrow\overline{{\mathbb{R}}^{n}}$ defined in some $D\in\mathfrak{D}$ and satisfying (1.3)–(1.4) at any point $x_{0}\in\overline{\mathfrak{D}},$ $0<r_{1}<r_{2}<\infty,$ for which there exists a continuum $A=A_{f}\subset D$ with $h(f(A))\geqslant\delta.$
@@
as $\varepsilon\rightarrow 0,$ where $A(x_{0},\varepsilon,\varepsilon_{0})$ is defined in (1.2). (If $x_{0}=\infty,$ (2.2) means the similar relation for $x_{0}=0$ with $\widetilde{Q}(x)=Q\left(\frac{x}{|x|^{2}}\right)$ instead of $Q(x)$).
@@
Proof.   We will mainly use the methodology that has been used for mappings of a fixed domain, see, e.g., [SevSkv]. Assume the contrary. Then there exists $a>0$ such that for $\delta=1/k,$ $k=1,2,\ldots,$ there are $m_{k}\in{\mathbb{N}}$ and $x_{k},x^{\,\prime}_{k}\in B(x_{0},1/k)\cap D_{m_{k}}$ such that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a.$ We may consider that the sequence $m_{k}$ is increasing by $k.$ Otherwise, $m_{k}=k_{0}$ for sufficiently large $k\in{\mathbb{N}}$ and some $k_{0}\in{\mathbb{N}},$ besides that, $x_{0}\in\overline{D_{k_{0}}}$ because $f_{k_{0}}$ is defined at $x_{k}\cap D_{k_{0}}$ and $x_{k}\rightarrow x_{0}$ as $k\rightarrow\infty.$ Now, from the assumption above we have that $h(f_{k_{0}}(x_{k}),f_{k_{0}}(x^{\,\prime}_{k}))\geqslant a.$ The latter contradicts with Lemma 2. So, $m_{k}$ may be chosen as an increasing sequence. Without loss of generality, going to renumbering, if required, we may consider that $h(f_{m_{k}}(x_{k}),f_{m_{k}}(x^{\,\prime}_{k}))\geqslant a$ holds for $m$ instead of $m_{k},$ i.e.,
@@
where $h(C_{m})$ denotes the chordal diameter of the set $C_{m}.$
Let $\Gamma_{m}$ be a family of paths joining $|\gamma_{m}|$ and $A_{m}$ in $D_{m},$ where $A_{m}$ is a continuum from the condition 3). By the definition of the class $\mathfrak{H}^{p}_{Q,\delta}(\mathfrak{D}),$ the relations (1.3)–(1.4) hold at $x_{0}.$ Therefore,
$M_{p}(f_{m}(\Gamma_{m}))\leqslant\int\limits_{A(x_{0},2^{\,-m},\varepsilon_{0})}Q(x)\cdot\eta^{p}(|x-x_{0}|)\ dm(x)$ (2.5)
for every Lebesgue measurable function $\eta:(2^{\,-m},\varepsilon_{0})\rightarrow[0,\infty]$ such that $\int\limits_{2^{\,-m}}^{\varepsilon_{0}}\eta(r)\,dr\geqslant 1.$ Observe that, the function
@@
On the other hand, observe that $f_{m}(\Gamma_{m})=\Gamma(C_{m},f_{m}(A_{m}),D_{m}^{\,\prime}).$ By the hypothesis of the lemma, $h(f_{m}(A_{m}))\geqslant\delta$ for all $m\in{\mathbb{N}}.$ Therefore, due to (2.4), $h(f_{m}(A_{m}))\geqslant\delta_{1}$ and $h(C_{m})\geqslant\delta_{1},$ where $\delta_{1}:=\min\{\delta,a/2\}.$ Using the fact that the domains $D_{m}^{\,\prime}:=f_{m}(D_{m})$ are equi-uniform with respect to $p$-modulus, we conclude that there exists $\sigma>0$ such that
@@
The following statement may be found in [Sev1, Lemmas 1.3, 1.4].
@@
$\int\limits_{\delta}^{\delta_{0}}\frac{dt}{t^{\frac{n-1}{p-1}}q_{x_{0}}^{\frac{1}{p-1}}(t)}<\infty,\qquad 0<\delta<\delta_{0},\qquad\int\limits_{0}^{\delta_{0}}\frac{dt}{t^{\frac{n-1}{p-1}}q_{x_{0}}^{\frac{1}{p-1}}(t)}=\infty\,.$
Then there exist a number $\varepsilon_{0}\in(0,1)$ and a function $\psi\geqslant 0$ satisfying (2.1) such that the relation (2.2) holds.
@@ 3 On image of the continuum under open discrete and closed mappings
Proposition 3.2.   Let $f:D\rightarrow{\mathbb{R}}^{n}$ be a discrete open and closed (boundary preserving) mapping, $\beta:[a,\,b)\rightarrow f(D)$ be a path, and $x\in\,f^{\,-1}\left(\beta(a)\right).$ Then $\beta$ has a total $f$-lifting $\alpha:[a,b)\rightarrow D$ starting at $x$ (see, e.g., [Vu, Lemma 3.7]). Moreover, if $\beta(t)\rightarrow\partial f(D)$ as $t\rightarrow b,$ then $\alpha(t)\rightarrow\partial D$ as $t\rightarrow b$ (see, e.g., Lemma 3.12 in [MRV]).
@@
Lemma 3.2.  Let $D_{0}\subset{\mathbb{R}}^{n},$ be a domain in ${\mathbb{R}}^{n},$ $n\geqslant 2,$ and let $\mathfrak{D}=\{D_{m}\}_{m=1}^{\infty}$ be a sequence of domains the kernel of which is $D_{0}.$ Let $A$ be a non-degenerate continuum in $D_{0},$ $A\subset\bigcap\limits_{m=1}^{\infty}D_{m},$ and let $Q:{\mathbb{R}}^{n}\rightarrow[0,\infty]$ be a Lebesgue measurable function. In addition, let $f_{m}:D_{m}\rightarrow{\mathbb{R}}^{n}$ be a sequence of open discrete and closed mappings satisfying relations (1.3)–(1.4) at any point $x_{0}\in\overline{D_{0}}$ for all $0<r_{1}<r_{2}<d(x_{0},\partial D_{m})$ such that $h(f_{m}(A))\geqslant\delta>0$ for any $m=1,2,\ldots$ and $p=n$ for some $\delta>0.$ Assume that, $Q$ has a finite mean oscillation in $A,$ or the relation (1.5) holds for every $x_{0}\in A$ and some $\beta(x_{0})>0.$ If there is $r>0$ such that $h(E)\geqslant r$ whenever $E$ is any component of $\partial f_{m}(D_{m}),$ then there exists $\delta_{1}>0$ such that $h(f_{m}(A),\partial f_{m}(D_{m}))\geqslant\delta_{1}$ for any $m=1,2,\ldots.$
Proof.   Observe that, $h(A,\partial D_{m})\geqslant r_{*}>0$ for some $r_{*}>0,$ and all $m\in{\mathbb{N}}$ except finite numbers. Otherwise, there are sequences $x_{m}\in A$ and $y_{m}\in\partial D_{m}$ with $h(x_{m},y_{m})\rightarrow 0$ as $m\rightarrow\infty.$ We may consider that $x_{m},y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty.$ Obviously, $x_{0}\in A$ by the closeness of $A.$ Now, there is $\varepsilon_{0}>0$ such that $\overline{B(x_{0},\varepsilon_{0})}\subset D_{0}$ and $\overline{B(x_{0},\varepsilon_{0})}$ belongs to all $D_{m}$ except of finite numbers. However, the latter contradicts the condition $y_{m}\rightarrow x_{0}$ as $m\rightarrow\infty$ while $y_{m}\in\partial D_{m},$ because in this case $\overline{B(x_{0},\varepsilon_{0})}\cap\partial D_{m}\neq\varnothing$ for infinitely many $m\in{\mathbb{N}}.$
@@
the relation (1.5) implies that, given $0<\widetilde{\varepsilon_{1}}<r_{*}\leqslant h(A,\partial D_{m})$ we may find $\varepsilon_{1}=\varepsilon_{1}(x_{0})$ such that $\int\limits_{\varepsilon_{1}}^{\widetilde{\varepsilon_{1}}}\frac{dt}{tq_{x_{0}}^{\,\prime\,\frac{1}{n-1}}(t)}>1\,.$ We cover $A$ with the balls $B(x_{0},\varepsilon_{1}),$ $x_{0}\in A,$ where $\varepsilon_{1}$ as above. By Heine-Borel-Lebesgue lemma we may find a finite subcover $B(x_{1},\varepsilon_{1}),$ $B(x_{2},\varepsilon_{2}),\ldots,B(x_{N},\varepsilon_{N})$ such that $A\subset\bigcup\limits_{i=1}^{n}B(x_{i},\varepsilon_{i})\,$, $A\subset\bigcup\limits_{i=1}^{n}B(x_{i},2\varepsilon_{i})\,.$ By the choice of $\varepsilon_{i}>0,$ $\varepsilon_{i}<\min\{\beta(x_{0}),\frac{1}{2}r_{*}\},$ we have that $B(x_{i},2\varepsilon_{i})\subset D_{m}.$ Since $D_{m}$ is a domain in ${\mathbb{R}}^{n},$ $m=1,2,\ldots,$ $\partial D_{m}\neq\varnothing.$ In addition, since $C(\partial D_{m},f_{m})\subset\partial f_{m}(D_{m})$ whenever $f_{m}$ is open, discrete and closed (see [Vu, Theorem 3.3]), $\partial f_{m}(D_{m})\neq\varnothing,$ as well. Thus, the quantity $h(f_{m}(A),\partial f_{m}(D_{m}))$ is well-defined, so the formulation of the lemma is correct.
@@
Put $P>0$ and $U=B_{h}(y_{0},r_{0})=\{y\in\overline{{\mathbb{R}}^{n}}:h(y,y_{0})<r_{0}\},$ where $2r_{0}:=\min\{r/2,\delta/2\},$ where $r$ and $\delta$ are numbers from the condition of the lemma. Observe that $E_{m}\cap U\neq\varnothing\neq E_{m}\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $y_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $y_{m}\in E_{m};$ besides that, $h(E_{m})\geqslant r>r/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Observe that, $E_{m}$ is closed (see [Ku, Theorem 1.III.5.46]). Now, $E_{m}$ is a continuum, so that $E_{m}\cap\partial U\neq\varnothing$ by Proposition 3. Similarly, $f_{m}(A)\cap U\neq\varnothing\neq f_{m}(A)\setminus U$ for sufficiently large $m\in{\mathbb{N}},$ since $x_{m}\rightarrow y_{0}$ as $m\rightarrow\infty,$ $x_{m}\in f_{m}(A);$ besides that $h(f_{m}(A))\geqslant\delta>\delta/2\geqslant 2r_{0}$ and $h(U)\leqslant 2r_{0}.$ Since $f_{m}(A)$ is a continuum, $f_{m}(A)\cap\partial U\neq\varnothing$ by Proposition 3. By the proving above, $E_{m}\cap\partial U\neq\varnothing\neq f_{m}(A)\cap\partial U$ for sufficiently large $m\in{\mathbb{N}}.$ By Lemma 3 there is $\subset U, VV$ is a neighborhood of $y_{0},$ such that
@@
for any continua $E,F\subset\overline{{\mathbb{R}}^{n}}$ with $E\cap\partial U\neq\varnothing\neq E\cap\partial V$ and $F\cap\partial U\neq\varnothing\neq F\cap\partial V.$ Arguing similarly to above, we may prove that $E_{m}\cap\partial V\neq\varnothing\neq f_{m}(A)\cap\partial V$ for sufficiently large $m\in{\mathbb{N}}.$ Thus, by (3.1)
@@
Let $\Delta_{m}$ be a family of all total $f_{m}$-liftings of $\Gamma_{m}$ starting at $A$ (it exists by Proposition 3). Now
@@
Let $\beta_{m}:[0,t_{m})\rightarrow D_{m},$ $\beta_{m}\in\Delta_{m},$ $f_{m}\circ\beta_{m}=\alpha_{m}.$ By Proposition 3 $\beta_{m}(t)\rightarrow\partial D_{m}$ as $t\rightarrow t_{m}.$
@@
Indeed, let $\gamma\in\Delta_{m},$ in other words, $\gamma:[0,t_{m})\rightarrow D_{m},$ $\gamma(0)\in A$ and $\gamma(t)\rightarrow\partial D_{m}$ as $t\rightarrow t_{m}-0.$ Now, there is $1\leqslant i\leqslant N$ such that $|\gamma|\cap B(x_{i},\varepsilon_{i})\neq\varnothing\neq|\gamma|\cap(D_{m}\setminus B(x_{i},\varepsilon_{i})).$ Therefore, by Proposition 3 there is $0<t^{\,*}_{1}<t_{m}$ such that $\gamma(t^{\,*}_{1})\in S(x_{i},\varepsilon_{i}).$ We may assume that $\gamma(t)\not\in B(x_{i},\varepsilon_{i})$ for $t>t^{\,*}_{1}.$ Put $\gamma_{1}:=\gamma|_{[t^{\,*}_{1},t_{m}]}.$ Similarly, $|\gamma_{1}|\cap B(x_{i},2\varepsilon_{i})\neq\varnothing\neq|\gamma_{1}|\cap(D_{m}\setminus B(x_{i},2\varepsilon_{i})).$ By Proposition 3 there is $t^{\,*}_{1}<t^{\,*}_{2}<t_{m}$ with $\gamma(t^{\,*}_{2})\in S(x_{i},2\varepsilon_{i}).$ We may assume that $\gamma(t)\in B(x_{i},2\varepsilon_{i})$ for $t<t^{\,*}_{2}.$ Put $\gamma_{2}:=\gamma|_{[t^{\,*}_{1},t^{\,*}_{2}]}.$ Then, the path $\gamma_{2}$ is a subpath of $\gamma,$ which belongs to the family $\Gamma(S(x_{i},\varepsilon_{i}),S(x_{i},2\varepsilon_{i}),A(x_{i},\varepsilon_{i},2\varepsilon_{i})).$ Thus, the relation (3.5) is established.